The measure of angle SQR is 2/5 the measure of angle PQR is 2/3 the measure of a right angle, what is the measure of a right angle, what is the measure of angle SQR.

HERE'S ANOTHER QUESTION
In Isosceles triangle ABC, the measure of angle A is 80 degrees. If another angle of the triangle measure x degrees, x ≠ 80, what is one possible value of x?

Here's Another
If r is directly proportional to s and if s= 2/3 when r=4/5, what is the value of r when s=4/9 ?

Answer by rfer(16322)

(2/3)/(4/5)=(4/9)/(r)
2/3r=16/45
r=(16/45)/(2/3)
r=16/45*3/2
r=48/90
r=8/15